OpenStudy (anonymous):

Hi Openstudy, I want to solve two coefficients: 2A + 2B = 6 and 2A + 2B = 4 i know i am usually suppose to subtract them but this is not linearly independent. My friend suggested that I substitute A to equal 1, so my equation will become 1 = (6-2B)/2 which would lead B to equal 2. My brother told me that this question is unsolvable, but i need to be able to solve this for Lando's equation. Can anyone please tell me which is true?

OpenStudy (agent0smith):

There is no solution. Think about it this way: how can (2A+2B) be equal to two different numbers? That would mean that 6 = 4.

OpenStudy (anonymous):

my other friend said that same exact thing

OpenStudy (anonymous):

hm what if i solve for A so it becomes A = (6-2B) / 2 then substitute it in the second problem?

OpenStudy (agent0smith):

Even if you use the substitution method, you won't get B = 2, you'll get 6 = 4. I think you made a mistake in your substitution. A = (6-2B)/2 = 3 - B put that into your other equation, and you'll still get 6 = 4. 2(3-B) + 2B = 4 6 - 2B + 2B = 4 6 = 4

OpenStudy (anonymous):

ic..

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

hey while we are on the topic, i need this answer to solve for the "2nd order homogenous linear recurrence relation with constant coefficients", you don't happen to know what I should do if this occurs do you?

OpenStudy (agent0smith):

hmm, sorry I can't help with that, too tired to look into it right now. This might help though? http://people.uncw.edu/tompkinsj/133/recursion/homogeneous.htm

OpenStudy (anonymous):

ill look into it, thanks a lot!